40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10105     Entry time: Wed Jun 25 20:45:04 2014
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: AS55 Bodeplot 


I finally did carry out a measurement on the network analyzer. This proves that the previous system will work properly. Just the optical splitter problem is to be taken care of.

For this, after Elog 10102, I did not touch any of the tables or photodiodes. Only turned on the laser at 1Y1 and took readings from the NA located nearby. I switched off the laser after measurements. The power to REF PD remains on.

I plotted transimpedence plots in the usual way and got ridiculous values of 15 ohms at 55MHz. Obviously there is the problem of varying amount of power illuminating the REF PD and AS55.

So, I just plotted the bode plots of transfer function got from the NA to check if the characteristics of AS55 looks as it was supposed to be and Yes! I got a nice peak at 55MHz.


Acquiring data


  • Diode Laser driving current: 240mA 
  • The following conditions were set on Network Analyzer Agilent 4395:


1) Frequency sweep range: 1MHz to 100 MHz.

2) Number of Points sampled in  the range: 801

3) Type of sweep: Linear

  • Set the NA to give the corresponding transfer function value (output of AS55 over output of 1611) and also Phase response in degrees.
  • Save the data into floppy disk for processing on the computer.



The experimental values obtained were:

DC output voltage of REF PD: 7.48V

DC output voltage of AS55: 53.7mV

Power incident on REF PD and AS55 respectively: 2.4mW  and 1.6mW

Taking the DC transimpedence of AS55 as 66.2 ohms (from schematic given at D1300586-v1) and for REF PD as 1E04 ohms

Hence, Responsivity for REF PD and AS55 respectively are:  0.312 A/W and 0.51 A/W


The data and code used are in the zip file.

Attachment 1: AS55_06-25-2014.zip  15 kB
Attachment 2: AS55_bodeplot.pdf  9 kB  | Hide | Hide all
ELOG V3.1.3-